# Coupled Flow-Thermal Analysis of a Rocket Nozzle with Charring Ablative Thermal Protection System

### v1

@@ Abstract
This study investigates conjugate heat transfer analysis of a rocket nozzle featuring a charring ablative thermal protection system. The study is carried out by weakly coupling a commercial computational fluid dynamics (CFD) flow solver Fluent with an in-house material thermal response solver. The coupling of the two solvers is carried out by exchanging the boundary conditions at the fluid-solid interface using a non-iterative approach. Validation of the numerical framework is performed. A blowing correlation is used to model the pyrolysis gas behavior at the interface. Results highlight a significant surface heating and ablation particularly at the nozzle throat.
@@ 3 Validation Case
In this study, ANSYS Fluent is utilized for the CFD simulations, while an in-house material thermal response solver is used for obtaining thermal response of a TPS material. The latter is already validated comprehensively in previous works [32, 35]. To validate the flow solver, ANSYS Fluent, a test case from the NASA Technical Report [30] is employed. The mesh of the flow domain is illustrated in Fig. 1. Details of the nozzle geometry are provided in Table 3. A structured mesh with 250 $\times$ 96 elements was used.
To accurately capture wall properties, the first layer of cells was positioned close to the wall to ensure adequate resolution of the boundary layer. Additionally, the mesh was refined near the nozzle wall to capture heat transfer phenomena effectively, maintaining the Y+ value below 1 throughout the nozzle wall. The test case parameters and boundary conditions are summarized in Table 3. An isothermal boundary condition was applied to the nozzle wall, with the wall temperature set to half of the nozzle inlet temperature.
@@
Nozzle inlet density 2.1306 kg/m3
@@
Figure 1: Nozzle with cooling duct mesh
@@
The flow properties like mach number Fig. 3a, 3b, temperature Fig. 4a, 4b and pressure Fig. 5a, 5b with their respective axial variation can be visualized in the plots provided.
@@ 4 Results And Discussion
After successful validation of flow solver, complete conjugate heat transfer analysis is performed. This time a thermal protection system (TPS) is introduced on the nozzle wall which is analyzed for the solid domain. TPS is made of AVCOAT 5026–39/HC-G, a glass-filled epoxy novolac ablative material. The thickness of 10 mm is provided for the solid domain simulations. AVCOAT material has an ablation temperature of 922 K and the relative thermo-physical properties of the material are taken from Ref. [36]. The mesh used for the flow simulations is same as used for validation case study, shown in Sec. 3. The mesh of the solid domain, as shown in Fig. 6, is having 250 cells in transverse direction and 50 cells in radial direction. The simulation results are provided in different time intervals in order to give a clear picture of results.
@@ 4.1 Comparison of Heat Flux and Surface Temperature
(a) Heat flux variation
(b) Surface temperature variation
Figure 7: Comparison of surface heat flux and surface temperature; Time interval: 0.2-1 s.
Figure 8a depicts heat flux variation during the time interval of one to five seconds. It can be observed that the heat flux generally shows a decreasing trend over time. However, during the 5th second, a reversal in this trend is noted. As seen in Fig. 8b, it can be noticed that the surface temperature of the TPS material near throat region reached the ablation temperature during the 5th second. The rise of surface temperature to this limit shows the material in the throat region goes under ablation earlier as compared to other regions.
(a) Heat flux variation
(b) Surface temperature variation
Figure 8: Comparison of surface heat flux and surface temperature; Time interval: 1-5 s.
@@
(a) Heat flux variation
(b) Surface temperature variation
Figure 9: Comparison of heat flux and surface temperature; Time interval: 5-10 s.
Between 10 and 15 s, the heat flux stabilizes across the nozzle wall. It becomes apparent that the increment in the heat flux is gradual, indicating it is approaching a steady state, as shown in Fig. 10a. From the temperature profile (Fig. 10b), it is clear that the most part of nozzle wall has undergone ablation during this time interval.
(a) Heat flux variation
(b) Surface temperature variation
Figure 10: Comparison of heat flux and surface temperature; Time interval: 10-15 s.
During the 15-20 s interval, the progression in the heat flux is minimal, as illustrated in Fig. 11a. The surface temperature of the nozzle wall, with the exception of entry section reached the ablation threshold of 922 K, except the entry part of the nozzle, as highlighted in Fig. 11b
(a) Heat flux variation
(b) Surface temperature variation
Figure 11: Comparison of heat flux and surface temperature; Time interval: 15-20 s.
@@
Simulation results were extended further in time, up to 120 s. Since the nozzle wall temperature remained at 922 K, no further increment in the heat flux was evident. The heat flux at 120 s matched completely with that at 22 s. At this point, the entire nozzle wall has undergone ablation. The temperature contour of the TPS material is presented in Fig. 4a.
(a) Heat flux variation
(b) Surface temperature variation
Figure 12: Comparison of heat flux and surface temperature; Time interval: 20-25 s
@@ 4.2 Throat Region Heat Transfer Analysis
The throat of a rocket nozzle is a critical region as it separates subsonic flow region from the supersonic flow, hence, a detailed analysis is needed. This region of the nozzle experiences maximum heat flux, as previously discussed. The results of heat flux variation with time for the throat region are highlighted in Fig. 13a. The results illustrate that, initially the heat flux is at its highest value, followed with a sudden decrease in the heat flux as time progresses. As demonstrated by the graph, it is clear that the ablation process begins at around 5 s. Once ablation starts, heat flux increases again for a few seconds. Further in time, one can see that heat flux attains a stable value before stabilizing at a constant value.
@@
Figure 14: Comparison of heat flux at three different X locations
@@
In comparison to the heat flux, the surface temperature of the throat region continues to increase over time, until it reaches the ablation temperature of 922 K, as shown in Fig. 13b. The temperature remains constant over time, once the surface of the throat reaches the ablation temperature. Simulation results are obtained up to 120 s, but results are shown only up to 60 s, since no further changes in the heat flux are observed beyond this time.
@@ 4.3 Comparative Heat Flux Analysis Across Nozzle Regions
To further understand the progression of heat transfer within the TPS, three locations are investigated, viz., the throat (X = 0.091 m), the converging section (X = 0.075 m), and the diverging section (X = 0.105 m) of the nozzle.The heat flux results are presented in Fig. 14. All three locations follow a similar trend, as discussed for the throat region. It can be noted that the maximum heat flux is noticed in the throat region, followed by the diverging section, with the converging section experiencing the least heat flux.
The temperature variation through the thickness of the TPS material at these locations is also analyzed, as depicted respectively in Figs. 15b, 15a and 16b. The surface near the high temperature gas flow reaches a notably higher temperature, while the back-end of the material remains at the adiabatic wall temperature. In addition to that, a gradual decrease in temperature through the thickness of the TPS material is observed.
(a) Temperature through the material at X= 0.091 m (throat)
(b) Temperature through the material at X= 0.075 m
Figure 15: Temperature variation at different locations.
(a) Temperature through the material at X= 0.105 m
(b) Temperature contour of TPS
Figure 16: Temperature through the material at X= 0.105 m and temperature contour of TPS.
@@ 4.4 Surface Recession
(a) Initial vs receded nozzle wall
(b) Total surface recession
@@ 5 Conclusions
A conjugate heat transfer analysis of rocket nozzle equipped with charring ablative thermal protection system was evaluated under the extreme heat loads. The effect of pyrolysis gas at the fluid and solid interface was modeled using blowing correction correlation. The coupled study of flow (steady state analysis) and transient thermal solver CATS enabled a detailed analysis of TPS of a rocket nozzle. The validation of the flow solver with the established experimental data demonstrated the reliability of our flow simulations. The thermal response modeled by in-house solver helped to understand the mechanisms, such as, ablation, charring and pyrolysis within the TPS material. Surface recession was calculated for 120 s, illustrating how the nozzle surface erodes away under extreme gas temperatures.
The maximum surface recession occurred in the throat region with a magnitude of approximately 2.5 mm at 120 s. These findings of ablation and surface recession underscore the need for enhanced heat insulation in such extreme heat load regions. The used approach presents a framework for predicting the thermal behavior of AVCOAT TPS material under extreme conditions. Nevertheless, there are areas for improvement. Future studies will be performed to understand the impact of coupling time-step better. Furthermore, this study did not incorporate mesh updates for the flow domain; incorporating mesh adaptation in future studies could capture the dynamics of flow better. This would enable even more accurate predictions of TPS performance and durability.
@@ Acknowledgments
The authors acknowledge the National Supercomputing Mission (NSM) for providing computing resources of “PARAM Sanganak” at IIT Kanpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and the Department of Science and Technology (DST), Government of India.

### v2

@@ Abstract
This paper presents a conjugate flow–thermal analysis of a rocket nozzle protected by a charring ablative thermal protection system (TPS). The study employs a coupled approach, integrating a CFD solver with an in-house transient material response code through the exchange of boundary conditions at the fluid–solid interface. The nozzle incorporates an AVCOAT TPS and is subjected to high-temperature compressible flow. Results identify the nozzle throat as the critical location, exhibiting the highest convective loading, early attainment of the material ablation temperature, and progressive surface recession. Temporal analysis of the coupled simulations reveals an initial peak in wall heat flux followed by a transient reduction and a subsequent resurgence as viscous dissipation and evolving surface conditions modify the near-wall thermal field. At 120s120~s of simulated operation, the maximum surface recession at the throat is approximately 2.5mm2.5~$m$. This research provides a methodology for predicting the thermal and ablative response of rocket nozzles equipped with charring TPS materials. The proposed framework offers valuable insights into the design and optimization of high-performance nozzles for extreme environments.
@@ 3 Validation Case
In this study, ANSYS Fluent is employed to perform the CFD simulations, while an in-house thermal response solver is used to predict the material response of the TPS. The latter has already been comprehensively validated in previous works [32, 35].
To validate the flow solver (ANSYS Fluent), a benchmark test case from the NASA Technical Report by Back et al. [30] is considered. Specifically, Case 262 is reproduced, which involves supersonic flow through a convergent–divergent nozzle with wall cooling. The configuration provides detailed experimental data on wall heat transfer and pressure distribution, making it a standard validation case for nozzle heat-transfer studies. The nozzle geometry and operating conditions are replicated from the experimental setup, and the computational mesh of the flow domain is shown in Fig. 1. Details of the nozzle geometry are provided in Table 3.
Figure 1: Nozzle with cooling duct mesh
@@
Nozzle inlet density 2.13 kg/m3
@@
The grid independence study was conducted to verify that the computed results were not influenced by mesh resolution. Three grids: coarse ($201\times 81$), medium ($251\times 96$), and fine ($301\times 101$) were tested. Figure 2a shows the variation of the heat transfer coefficient, where minimal differences were observed between the medium and fine grids. To accurately capture wall properties, the first layer of cells was positioned close to the wall to ensure adequate resolution of the boundary layer. Additionally, the mesh was refined near the nozzle wall to capture heat transfer phenomena effectively, maintaining the Y+ value way below 1 throughout the nozzle wall. Figure 2b confirms that the $Y^{+}$ values remained below 1 for all grids, ensuring accurate resolution of the viscous sublayer. A structured mesh with 251 $\times$ 96 elements was used for our simulations.
(a) Heat transfer coefficient for different grid resolutions.
(b) $Y^{+}$ distribution along the nozzle wall for the medium grid.
Figure 2: Grid independence and near-wall resolution assessment for the nozzle flow simulations.
The test case parameters and boundary conditions are summarized in Table 3. An isothermal boundary condition was applied to the nozzle wall, with the wall temperature set to half of the nozzle inlet temperature.
@@
The flow properties like mach number Fig. 4a, 4b, temperature Fig. 5a, 5b and pressure Fig. 6a, 6b with their respective axial variation can be visualized in the plots provided.
@@ 4 Results And Discussion
After successful validation of flow solver, complete conjugate heat transfer analysis is performed. This time a thermal protection system (TPS) is introduced on the nozzle wall which is analyzed for the solid domain. TPS is made of AVCOAT 5026–39/HC-G, a glass-filled epoxy novolac ablative material. The thickness of 10 mm is provided for the solid domain simulations. AVCOAT material has an ablation temperature of 922 K and the relative thermo-physical properties of the material are taken from Ref. [36]. The mesh used for the flow simulations is same as used for validation case study, shown in Sec. 3. The mesh of the solid domain, as shown in Fig. 7, is having 250 cells in transverse direction and 50 cells in radial direction. The boundary conditions for the further study are summarized in Table 4. Initially, an isothermal boundary condition was imposed on the nozzle wall, with the wall temperature set to half the nozzle inlet temperature at the initial time step. As the simulation progressed, a user-defined function (UDF) was employed to apply a spatially varying temperature distribution on the nozzle wall.
Table 4: Boundary conditions and operating parameters for the nozzle with Thermal Protection System (TPS).
Condition Value
Nozzle inlet pressure 344048.40 N/m2
Nozzle inlet temperature 1111.11 K
Nozzle inlet density 1.07 kg/m3
Transport model Sutherland’s law
Turbulence model BSL
Exit Mach number 2.5
Simulation results were evaluated at various time intervals to capture the thermal response of the nozzle with the TPS layer, providing insights into the material’s performance under operational conditions.
@@ 4.1 Comparison of Heat Flux and Surface Temperature
(a) Heat flux variation from 0.2–1.0 s
(b) Surface temperature at 1.0 s
Figure 8: Comparison of surface heat flux (0.2–1.0 s) and surface temperature (1.0 s).
Figure 9a shows the temporal evolution of the wall heat flux at the nozzle throat over the interval 1–5 s. The heat flux decreases during the early portion of this interval but reverses and increases during the fifth second. This behavior is examined in detail later in Sec. 4.2 and is attributed to an increase in near-wall viscous dissipation together with changes in the TPS surface temperature. As shown in Fig. 9b, the TPS surface temperature at the throat attains the ablation threshold (922 K) during the fifth second, indicating that ablation initiates earlier at the throat than at other axial locations. The onset of ablation and the resulting modification of the surface boundary condition steepen the local wall-normal temperature gradient, which explains the observed rise in instantaneous wall heat flux.
(a) Heat flux variation from 1.0-5.0 s
(b) Surface temperature variation at 5.0 s
Figure 9: Comparison of surface heat flux (1.0-5.0 s) and surface temperature (5.0 s).
@@
(a) Heat flux variation from 5.0-10.0 s
(b) Surface temperature variation at 10.0 s
Figure 10: Comparison of heat flux (5.0-10.0 s) and surface temperature (10.0 s).
Between 11 and 15 s, the heat flux stabilizes across the nozzle wall. It becomes apparent that the increment in the heat flux is gradual, indicating it is approaching a steady state, as shown in Fig. 11a. From the temperature profile (Fig. 11b), it is clear that the most part of nozzle wall has undergone ablation during this time interval.
(a) Heat flux variation from 11.0-15.0 s
(b) Surface temperature variation at 15.0 s
Figure 11: Comparison of heat flux (11.0-15.0 s) and surface temperature (15 s).
During the 16-20 s interval, the progression in the heat flux is minimal, as illustrated in Fig. 12a. The surface temperature of the nozzle wall, with the exception of entry section reached the ablation threshold of 922 K, except the entry part of the nozzle, as highlighted in Fig. 12b
(a) Heat flux variation from 16.0-20.0 s
(b) Surface temperature variation at 20.0 s
Figure 12: Comparison of heat flux (16.0-20.0 s) and surface temperature (20.0 s).
@@
(a) Heat flux variation from 20.0-25.0 s
(b) Surface temperature variation at 25.0 s
Figure 13: Comparison of heat flux (20.0-25.0 s) and surface temperature (25 s)
Simulation results were extended further in time, up to 120 s. Since the nozzle wall temperature remained at 922 K, no further increment in the heat flux was evident. The heat flux at 120 s matched completely with that at 22 s. At this point, the entire nozzle wall has undergone ablation.
@@ 4.2 Throat Region Heat Transfer Analysis
The throat of a rocket nozzle is a critical region as it separates the subsonic flow region from the supersonic flow, hence a detailed analysis is needed. This region of the nozzle experiences maximum heat flux, as discussed previously. The results of heat flux variation with time for the throat region are highlighted in Fig. 14a. The results illustrate that initially the heat flux is at its highest value, followed by a sudden decrease in the heat flux as time progresses. As demonstrated in the graph, the ablation process begins at around 5 s, at which point the heat flux increases again for a few seconds. Further in time, one can see that the heat flux attains a stable value.
@@
In comparison to the heat flux, the surface temperature at the throat continues to increase with time until it reaches the ablation temperature of 922 K, as shown in Fig. 14b. Once the throat surface attains the ablation temperature, the temperature remains fixed thereafter. Simulations were carried out up to 120 s, but results are displayed only up to 60 s for heat flux and 10 s for surface temperature, since no further changes in these quantities were observed beyond those times.
4.3 Thermal Field Characteristics
To interpret the observed heat flux trends, it is first necessary to examine the local flow properties near the nozzle throat. For this purpose, the static temperature was evaluated at the throat to analyze its radial variation from the nozzle centerline ($y=0$) to the cooled wall.
Figure 15a shows the radial static temperature distribution over the interval 3–6 s. This time interval was chosen to capture the flow properties before and after ablation initiates. The temperature profile exhibits three distinct regions: an initial gradual decrease, a local rise, and a final drop to the wall temperature.
In the outer portion of the boundary layer, mixing with slower and slightly cooler fluid, together with heat transfer toward the cooled wall, produces a modest reduction in temperature relative to the core ($y=0$). Closer to the wall, within the inner boundary layer, the steep reduction in axial velocity generates strong velocity gradients. The associated viscous dissipation converts kinetic energy into internal energy, producing a local recovery effect that exceeds the immediate conductive losses. This mechanism accounts for the observed rise in the temperature profile. The gas temperature is constrained by the prescribed cold-wall boundary condition; since the wall temperature is maintained below the recovery temperature, the static temperature falls sharply, reaching the prescribed surface temperature.
The overall sequence of decrease, local increase, and final decrease in temperature therefore reflects the interplay of three competing mechanisms: advective and mixing-induced cooling, viscous dissipation (recovery heating), and conductive removal into the wall. At $t=4.6$ s the throat surface reaches the ablation temperature and is thereafter held at 922 K. With increasing time, viscous dissipation progressively elevates the near-wall gas temperature, steepening the local wall-normal gradient $(\partial T/\partial y)_{w}$. As a result, the instantaneous wall heat flux $q_{w}$ initially increases before gradually decaying as the system relaxes toward a new quasi-steady state.
(a) Radial distribution of static temperature at the nozzle throat.
(b) Temperature gradients at the nozzle throat.
Figure 15: Static temperature and temperature gradients across the nozzle throat at different times.
Figure 15b presents the temperature gradients at the nozzle throat for times between 3 s and 6 s. These trends are consistent with the corresponding static temperature profiles and the observed heat flux variation. A distinct sign change in the gradient curves indicates the presence of a local temperature overshoot near the wall. As this overshoot strengthens with time, the near-wall gradients become sharper, leading to a more negative slope. This behaviour reflects the dynamic balance between conduction into the cooled wall and viscous dissipation within the boundary layer.
At $t=3.0$ s, the wall heat flux is highest owing to the large wall–fluid temperature difference. By $t=4.0$ s the gradient magnitude decreases and the heat flux correspondingly weakens. However, at later times (5–6 s) viscous heating remains significant and because the wall is held at a constant temperature of 922 K during this interval, the near-wall overshoot is enhanced, restoring a steeper gradient. Consequently, the wall–fluid temperature contrast increases again, producing a rise in the instantaneous wall heat flux. Overall, the temporal evolution of the heat flux can be understood as the outcome of competing mechanisms: thermal conduction into the cold wall versus viscous heating within the flow.
Figure 16: Radial distribution of total temperature at the nozzle throat.
Figure  16 illustrates the total temperature distribution across the nozzle throat for times between 3 s and 6 s. In the core flow, the total temperature remains nearly uniform and constant, consistent with the freestream value. However, within the near-wall region, a gradual increase in total temperature is observed over time. At $t=3.0$ s, the cooled wall extracts thermal energy efficiently, resulting in a lower near-wall total temperature. At later times (4–6 s), viscous dissipation within the boundary layer becomes more pronounced, leading to a slight but systematic rise in the total temperature near the wall. This behavior aligns with the static temperature overshoot and the steepening of temperature gradients previously discussed, confirming the competing influences of wall cooling and viscous heating in the throat region.
4.4 Density and Velocity Profiles at the Throat
Since density is governed by the ideal gas law, it strongly depends on the local static temperature. From the density profile, as shown in Fig.  17a, moving outward from the core flow toward the onset of the boundary layer, the density remains nearly constant. However, near the onset of the boundary layer, the temperature plot in Fig.  15a shows a noticeable drop in static temperature, and accordingly the density increases in this region. Further into the boundary layer, viscous dissipation causes the static temperature to rise sharply, which leads to a local dip in density. Finally, at the cold wall, the density increases again in accordance with the temperature trend dictated by the ideal gas law.
The velocity profile in Fig. 17b shows that axial velocity increases in the same region where the density decreases, consistent with the formation of a low-density pocket driven by viscous heating. To satisfy local mass conservation, the velocity overshoots in this zone so that the local mass flux is maintained, before decaying to zero at the wall due to the no-slip condition.
(a) Radial density profiles at the nozzle throat from 3 to 6 seconds.
(b) Axial velocity profiles at the nozzle throat from 3 to 6 seconds.
Figure 17: Density and axial velocity distributions across the nozzle throat at different times.
@@ 4.5 Comparative Heat Flux Analysis Across Nozzle Regions
To further elucidate the evolution of heat transfer in the TPS, three axial locations are examined: the throat (x=0.091mx=0.091\ $\mathrm${m}), the converging section (x=0.075mx=0.075\ $\mathrm${m}), and the diverging section (x=0.105mx=0.105\ $\mathrm${m}). Results are presented for the first 60 s of simulated time.
The heat flux results are presented in Fig. 18. All three locations follow a similar trend, as discussed for the throat region. The maximum heat flux occurs in the throat region, followed by the diverging section, while the converging section experiences the least heat flux.
The temperature variation through the thickness of the TPS material at these locations is analyzed in Figs. 19a, 19b, and 20a. The surface exposed to the hot gas flow reaches a significantly higher temperature, while the back-face remains close to the adiabatic wall temperature. A gradual decrease in temperature through the TPS thickness is observed. In addition, the overall thermal response of the nozzle wall is illustrated by the temperature contour of the TPS material after 120 s, shown in Fig. 20b.
Figure 18: Comparison of surface heat flux at three axial locations in the nozzle.
(a) Throat region (x=0.091mx=0.091\ $\mathrm${m})
(b) Converging section (x=0.075mx=0.075\ $\mathrm${m})
Figure 19: Temperature variation through the TPS thickness at the throat and converging section.
(a) Diverging section (x=0.105mx=0.105\ $\mathrm${m})
(b) Temperature contour of the TPS material after 120 s
Figure 20: Temperature variation at the diverging section and overall TPS temperature contour after 120 s.
@@ 4.6 Surface Recession
(a) Comparison of initial and receded nozzle wall geometries after 120 s of operation.
(b) Axial distribution of surface recession along the nozzle wall.
@@ 5 Conclusions
A conjugate flow–thermal study of a convergent–divergent rocket nozzle protected by a charring ablative TPS (AVCOAT) was presented. The work coupled a commercial CFD solver (ANSYS Fluent) with an in-house transient material response code (CATS) using a weak, non-iterative exchange of boundary conditions at the fluid–solid interface. Pyrolysis-gas effects were accounted for via the Kays [33] blowing correction. The flow solver was validated against the NASA benchmark by Back et al. (Case 262) [30], and the material solver was validated in prior studies [32, 35]. Together, these validations provide confidence in the coupled predictions.
The coupled simulations identified the nozzle throat as the most thermally critical location: the throat experiences the highest convective loading, attains the TPS ablation temperature earlier than adjacent regions, and exhibited the largest cumulative surface recession. Over 120 s of simulated operation the maximum surface recession at the throat is approximately 2.5 mm. Temporally, the wall heat flux exhibited an early peak, a transient reduction, and a later resurgence; this behaviour was explained by the competing effects of conduction to the cooled wall, viscous dissipation within the near-wall flow, and the evolving surface temperature as ablation and charring progressed. The model captured interacting mechanisms including pyrolysis, blowing corrections, viscous heating and conduction, thereby providing a predictive framework for TPS response under extreme nozzle heat loads. Overall, the presented framework offers a practical and physically consistent approach to predict AVCOAT TPS behaviour under extreme nozzle environments and highlights throat regions as priority locations for enhanced thermal protection and further study.
@@ Acknowledgments
The authors acknowledge the financial support provided by the Indian Space Research Organization (ISRO) through the Grant No. STC/AE/2024427L. The authors also thank the National Supercomputing Mission (NSM) for providing computing resources of “PARAM Sanganak” at IIT Kanpur, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and the Department of Science and Technology (DST), Government of India.